An electron travels at a speed of 5.11×10^7 m/s along the z-axis. What is the magnetic field at the point (0.00 cm,9.00 cm,9.00 cm) when the electron passes through the origin? -3.57×10-17 î T, 3.57×10^-17 î T, -7.14×10^-17 î T, 7.14×10^-17 î T,
Answer #1

position vector

r = 0.09j + 0.09 k

field at the given position due to moving charge

B = ( uo/ 4 pi) q ( v × r) / r^3

B = 10^-7* 1.6*10^-19* * 10^7*( 5.11 k × ( 0.09j + 0.09 k) ) / ( 0.09^2 + 0.09^2)^1.5

B = 7.76*10^-17* ( - 5.11* 0.09 i)

B = - 3.57*10^-17 i

option (a) is correct

========

Comment in case any doubt, will reply for sure.. Goodluck